I would like to know how to divide a negative number with a decimal.
example: - 64/0.8 = ?

Answer by stanbon(75887)

I would like to know how to divide a negative number with a decimal.
example:
Let x = -64/0.8
--------
x =-64/(8/10
---
x = -640/8
---
x = -80
=======
Cheers,
Stan H.
=======